OpenStudy (awkwardpanda):

2. what is the 50th term of the sequence that begins -6,0,6,12...? a. 292 b.288 c.286 d.290

OpenStudy (anonymous):

my guess is, since you are adding \(6\) each time, it is \[-6+499\times 6\]

OpenStudy (anonymous):

oops make that \[-6+49\times 6\]

OpenStudy (anonymous):

gosh for once it isn't C!

OpenStudy (awkwardpanda):

Thank you xD

OpenStudy (anonymous):

yw

OpenStudy (imstuck):

I personally recommend to those who struggle with this that they use the formula in order to keep things straight and be able to apply the formula to other problems. For an arithmetic sequence such as this the formula is \[a _{n}=a _{1}+(n-1)d\]where n is the number term you are looking for (here it is the 50th term), a1 is the first number in the sequence (here it is -6), and d is the difference between all the terms you are given (here it is 6). Fill in your formula. \[a _{50}=-6+(50-1)6\]It's exactly what satellite73 said, but it's just the formula you can use everytime (as long as its arithmetic and not geometric, for example).